A rectangle, 4.5 metres wide has the same area as a square with 12 metre sides, what are the dimensions of the rectangle?
please show exactly how its worked out, thank you

Answer by Murthy(22)

Let, L be the length of the rectangle and W be its width = 4.5 units
Let a be the length of the side of the square =12 units
Given, Area of rectangle = Area of square
L x W = 12^2
L= 144/4.5=32 units
Dimensions of rectangle are 4.5 and 32 units